# Quantifying the Effect of Test Set Contamination on Generative Evaluations

### v1

@@ 4 Further Training: Overtraining & Supervised Finetuning (SFT)
tokens per model, where $m$ is the overtraining multiplier and $N$ is the number of model parameters. We swept $m\in\{1,2,4,8,16\}$. Following Sardana et al. (2024); Gadre et al. (2024), we term $m=1$ “compute-optimal training” and term $m>1$ “overtraining”. Crucially, in Bordt et al. (2025) and here, as the overtraining multiplier increases, the additional tokens are new fresh non-repeated tokens; this differs from more practical settings where models might see select documents repeated tens-to-hundreds of times (Hernandez et al., 2022) or the entire corpus repeated for 4+ epochs (Muennighoff et al., 2023; Fang et al., 2025).
We find an interesting interaction between contamination and overtraining (Fig. 3): for models with low contamination, cross entropy on the MATH test set decreases with increasing overtraining, but for models with high contamination, cross entropy on the MATH set increases with overtraining. The cross-over point between test set replicas and overtraining multiplier shifts with model size: the crossover point falls from 32 test set replicas for 3434M parameter models to 10 replicas for 6363M parameter models to 1 replica for 9393M parameter models. Thus, as models become larger, the performance boost from contamination diminishes when overtraining with fresh data. Our interpretation is that while more fresh data is generally useful for improving model performance generally, it dilutes the “dosage” of the contaminated data, weakening how the model “responds” (as measured by task performance) (Schaeffer et al., 2025b).
Finding #4: Supervised Fine-Tuning on Training Set Has Opposing Effects, Depending on Pretraining Contamination
After pretraining, the first post-training step is oftentimes supervised finetuning (SFT) (Wei et al., 2022b; Ouyang et al., 2022). We turned to assessing what effect, if any, SFT has on contaminated pretrained models. Kocyigit (2025) recently studied this question and found that SFTing on the train set improves performance on the test set. However, as a key methodological difference, Kocyigit (2025) induced test set contamination via continued pretraining (Jin et al., 2022; Jang et al., 2022; Ibrahim et al., 2024; Parmar et al., 2024; Yildiz et al., 2025), whereas we introduced test set contamination uniformly throughout pretraining.
@@
One might expect that SFTing on the train set should increase test set performance across the board. We discover that, surprisingly, the opposite is sometimes true: SFTing on the train set can both help and hurt model performance on the test set, depending on the amount of contamination in pretraining (Fig. 4). For models with no or low $(<10)$ test set contamination, SFTing on the train set significantly reduces loss on the test set (purple). At $10$ test set replicas, SFT has no effect (aqua), but as contamination increases, SFTing on the train set significantly increases loss on the test set (yellow, green). We conjecture that training highly contaminated models on the MATH train set causes forgetting of the test set in tandem with improved generalization. As studied in Section 1, the benefits of contamination on test loss dwarf those of generalization (which would be improved by SFT on the train set). Therefore, the net impact of SFTing on the train set for highly contaminated models is counterintuitively to increase cross entropy loss.
5 Generation: Sampling Temperature and Solution Length
Generative benchmarks introduce distinct complexities compared to discriminative evaluations. Unlike multiple-choice tasks where predetermined sequences are score, generative solutions require the model to produce a coherent sequence over hundreds of tokens or more. This introduces two critical levers for analyzing contamination: sampling temperature and solution length. In this section, we demonstrate that while contamination mimics reasoning capability under greedy decoding, this competence is brittle. We show that sampling and long-context requirements act as barriers that memorization struggles to overcome.
@@ 5 Generation: Sampling Temperature and Solution Length
Finding #5: High Temperature Sampling Mitigates the Gains from Contamination
@@
We evaluated the pretrained models using temperature-only sampling, sweeping from 0 (“greedy”) to $1.5$. We observe that Math Verify scores remain stable between greedy decoding and low-temperature sampling ($\tau\leq 0.56$) (Fig. 5, left and center). However, as temperature increases beyond this point, performance degrades quickly.
Crucially, higher temperature sampling acts as an equalizer, significantly reducing the performance gap between highly contaminated and uncontaminated models. While uncontaminated models are bounded by a performance floor, contaminated models lose their memorization advantage under stochastic sampling. For example, increasing temperature from 0 to $1$ causes performance at high contamination levels ($1000$ replicas) to collapse by a factor of 40, bringing them much closer to the baseline performance of uncontaminated models. This suggests that contamination-driven performance is brittle; small adjustments in inference settings can eliminate the effects of memorization almost entirely. High-temperature sampling flattens these peaks, causing the model to diverge from the memorized path and exposing the “illusion” of competence created by leakage.
Figure 6: Performance Declines with Increasing Solution Token Length. Math Verify Scores decrease exponentially with solution length at high levels of benchmark replication. For lower levels of benchmark replication, score decreases appear to follow approximate power laws. Trends are consistent between the 153M and 344M parameter models. Shaded regions represent 95% confidence intervals.
Figure 7: Math Verify Score Is Correlated with Pretraining Loss. Math Verify scores correlate strongly with the cross entropy loss achieved on the MATH Test Set during training, where differences in these graphs are attributable to increased repetition on the benchmark test set. The correlation is significantly weaker for high temperatures, and falls to nearly 0 for temperatures above 11.
Finding #6: Longer Solutions Reveal a Shift from Power Law to Exponential Decay
@@
To understand how solution length constrains performance, we binned problems into 10 log-spaced intervals ranging from the shortest (15 tokens) to the longest (1949 tokens). We find that Math Verify scores decrease significantly as solution length increases (Fig 6). We identify a striking shift in the functional form of this decay based on contamination level. At lower contamination levels, the decay follows an approximate power law, consistent with robust reasoning capabilities. However, at the highest levels of contamination, this decay becomes exponential with respect to token length. This suggests that maintaining a coherent memorized chain becomes increasingly difficult as the sequence grows—a single error can cause the model to decohere from the exact memorized text. This aligns with recent findings that longer sequences require more repetitions to be memorized (Jiang et al., 2025; Lu et al., 2024), though we demonstrate this here through controlled pretraining.
@@
Furthermore, solution length interacts with sampling temperature. While increasing temperature hurts performance generally, it causes catastrophic failure for long solutions. For short solutions ($\leq 100$ tokens) at high contamination (316 replicas), raising the temperature from 0 to $1.0$ drops accuracy by $\sim 45\%$ on the largest model. However, for solutions of $400$ tokens, the same temperature increase causes accuracy to drop by nearly $100\%$. This highlights a key distinction between generative and discriminative evaluations: unlike multiple-choice tasks where temperature is largely irrelevant, in generative settings, the combination of solution length and stochastic sampling can completely reverse the gains from even extreme levels of memorization.
Finding #7: Generative Performance Closely Tracks Cross-Entropy Loss, but These Are Decoupled at High Sampling Temperatures
We find a strong negative correlation between a model’s Math Verify score and its cross-entropy loss on the MATH test set (Fig. 7). Under greedy or low-temperature sampling, decreasing test set loss (via increased contamination) translates directly into a sharp, non-linear increase in generative performance. However, higher temperatures decoupled these metrics. While highly contaminated models achieve low cross-entropy loss (indicating they assign high probability to the correct next token), sampling at high temperatures prevents them from staying on the narrow path of the exact solution. Consequently, at temperatures above $1.0$, the correlation between cross entropy loss and Math Verify score drops to nearly 0.
@@ 6 Discussion
Our application of neural scaling laws reveals the economic magnitude of this threat. We demonstrate that a single replica of the test set allows a model to bypass the “irreducible error” barrier, effectively simulating performance that would otherwise require an infinite amount of compute on uncontaminated data. However, this performance is brittle. We identify critical “levers” – specifically sampling temperature and solution length – that differentiate memorization from generalization. Unlike robust reasoning, which survives stochastic sampling, contamination-driven performance collapses under high temperatures. Similarly, the shift from power-law to exponential decay in performance as solution length increases signals the limits of the model’s ability to hold a memorized chain without decoherence.
Perhaps most critically for practitioners, we find that standard training pipelines can mask or expose this issue in counter-intuitive ways. The interaction between contamination and further training is governed by a “dosage” effect (Schaeffer et al., 2025b): overtraining on fresh data dilutes the contamination, and Supervised Fine-Tuning (SFT) on valid training data can actually degrade test performance by overwriting memorized test samples. This implies that a drop in test accuracy after SFT, usually a sign of alignment tax or forgetting, may actually be a positive signal of decontamination.
@@
We focused on a single generative benchmark, MATH, to enable precise automatic verification and controlled contamination. Consequently, our findings may not fully capture how contamination behaves for tasks with higher entropy, such as open-ended dialogue or creative writing. Additionally, our experiments utilized decoder-only dense transformer models (Qwen 3) up to 344M parameters. Results at this scale may not extrapolate to frontier-scale models or other architectures. Finally, our pretraining corpus represents a specific mixture of web-crawl data; specialized or heavily filtered corpora could alter the base difficulty of memorization and the specific thresholds at which contamination effects become visible.
@@
Our findings suggest several avenues for future research in generative evaluation:
  • Reconciling Memorization Thresholds: Our discovery that a single test set replica drives loss below the irreducible error contrasts with recent assertions that single-shot verbatim memorization is an “illusion” (Huang et al., 2024) or that membership inference is fundamentally capped (Hayes et al., 2025). Future work should investigate what the explanatory differences are.
  • Inference-Time Detection Methods: Since high-temperature sampling and long-context generation disproportionately harm contaminated models, future work could develop lightweight “stress tests” that sweep temperature or solution length to detect contamination without access to the pretraining corpus.
  • Architectural Susceptibility: We studied dense models, but it remains an open question whether Mixture-of-Experts (MoEs) or State Space Models (SSMs) exhibit different memorization capacities. Do MoEs isolate contaminated memories in specific experts, making them easier to excise?
  • Harder Benchmarks: We hypothesize that contamination efficacy is inversely correlated with problem compressibility. Investigating whether harder benchmarks (e.g., AIME) require higher “dosages” of leakage to achieve the same performance boost would be a valuable contribution to benchmark design.
@@
An earlier version of this manuscript was peer reviewed and presented at the NeurIPS 2025 Workshop on Evaluating the Evolving LLM Lifecycle: Benchmarks, Emergent Abilities, and Scaling (Schaeffer et al., 2025c).
@@ 7 Acknowledgments
RS acknowledges support from Stanford Data Science and from the OpenAI Superalignment Fast Grant. JK acknowledges support from NSF grant number DGE1656518. SK acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc. KZL is generously supported by the Amazon AI PhD Fellowship. AA is generously supported by a Knight-Hennessy Fellowship, an NSF Graduate Research Fellowship, and a Georgetown Foundation Research Grant.
@@ Appendix B Pretraining Implementation Details
We pretrained Qwen 3 (Yang et al., 2025a) architecture causal language models. We used the AdamW optimizer (Loshchilov & Hutter, 2019) with HuggingFace defaults. We used linear warmup for 250 steps followed by cosine annealing. Following Shuai et al. (2024), we set the number of tokens per optimizer step $\propto D^{0.264}$. The base learning rate was 1e61e-6 and the maximum learning rate was the base learning rate times the square of the number of tokens per optimizer step. Gradients were clipped to a maximum of 1.0. For more information, please see our public GitHub repository.

### v2

@@ 4 Further Training: Overtraining & Supervised Finetuning
tokens per model, where $m$ is the overtraining multiplier and $N$ is the number of model parameters. We swept $m\in\{1,2,4,8,16\}$. Following Sardana et al. (2024); Gadre et al. (2024), we term $m=1$ “compute-optimal training” and term $m>1$ “overtraining”. Crucially, in Bordt et al. (2025) and here, as the overtraining multiplier increases, the additional tokens are new, fresh, non-repeated tokens; this differs from more practical settings where models might see select documents repeated tens-to-hundreds of times (Hernandez et al., 2022) or the entire corpus repeated for 4+ epochs (Muennighoff et al., 2023; Fang et al., 2025).
Figure 4: Supervised Finetuning on the Train Set Has Opposing Effects, Depending on Pretraining Contamination. For models pretrained with little-to-no contamination (<10<10 test set replicas), supervised finetuning (SFT) on the MATH train set decreases loss on the test set. For models pretrained with more contamination (>10>10 test set replicas), SFT on the train set increases loss on the test set. We conjecture that during SFT, contaminated models learn to generalize but also forget their contaminated pretraining data, and the effects of contaminated test set data are more impactful than generalization for small models, leading to a net increase in test loss.
We find an interesting interaction between contamination and overtraining (Fig. 3): for models with low contamination, cross entropy on the MATH test set decreases with increasing overtraining, but for models with high contamination, cross entropy on the MATH set increases with overtraining. The cross-over point between test set replicas and overtraining multiplier shifts with model size: the crossover point falls from 32 test set replicas for 3434M parameter models to 10 replicas for 6363M parameter models to 1 replica for 9393M parameter models. Thus, as models become larger, the performance boost from contamination diminishes when overtraining with fresh data. Our interpretation is that while more fresh data is generally useful for improving model performance generally, it dilutes the “dose” of the contaminated data, weakening how the model “responds” (as measured by task performance) (Schaeffer et al., 2025c).
Finding #5: Supervised Fine-Tuning on Training Set Has Opposing Effects, Depending on Pretraining Contamination
After pretraining, the first post-training step is oftentimes supervised finetuning (SFT) (Wei et al., 2022b; Ouyang et al., 2022). We turned to assessing what effect, if any, SFT has on contaminated pretrained models. Kocyigit (2025) recently studied this question and found that SFT on the train set improves performance on the test set. However, as a key methodological difference, Kocyigit (2025) induced test set contamination via continued pretraining (Jin et al., 2022; Jang et al., 2022; Ibrahim et al., 2024; Parmar et al., 2024; Yildiz et al., 2025), whereas we introduced test set contamination uniformly throughout pretraining. The MATH format was preserved between pretraining and SFT.
@@
Figure 6: Performance Declines with Increasing Solution Token Length. Math Verify Scores decrease with solution length, with the effects modulated by temperature. At high rates of contamination and higher temperatures, math verify scores fall exponentially with the solution length. At lower temperatures, and high rates of contamination, solution length has almost no effect. Models exposed to less contamination also exhibit declining scores with longer solution lengths, but the dropoff is concave up and occurs much more slowly.
@@
One might expect that SFT on the train set should increase test set performance across the board. We discover that, surprisingly, the opposite is sometimes true: SFT on the train set can both help and hurt model performance on the test set, depending on the amount of contamination in pretraining (Fig. 4). For models with no or low $(<10)$ test set contamination, SFT on the train set significantly reduces loss on the test set (purple). At $10$ test set replicas, SFT has no effect (aqua), but as contamination increases, SFT on the train set significantly increases loss on the test set (yellow, green). We conjecture that during SFT, contaminated models learn to generalize but also forget their contaminated pretraining data, and the effects of contaminated test set data are more impactful than generalization for small models, leading to a net increase in test loss. As studied in Section 3, the benefits of contamination on test loss dwarf those of generalization (which would be improved by SFT on the train set). Therefore, the net impact of SFT on the train set for highly contaminated models is counterintuitively to increase cross entropy loss.
5 Inference: Temperature & Solution Length
Finally, we arrive at the last stage of the model lifecycle: inference. When models are deployed for generation, inference-time parameters—particularly sampling temperature and the length of generated solutions—introduce new considerations for how contamination manifests.
@@ 5 Inference: Temperature & Solution Length
Finding #6: High Temperature Sampling Mitigates Gains from Contamination
@@
We evaluated the pretrained models using temperature-only sampling, sweeping from 0 (“greedy”) to $1.5$. We observed that Math Verify scores remain stable between greedy decoding and low-temperature sampling ($\tau\leq 0.56$) (Fig. 5, left and center). However, as temperature increases beyond this point, performance degrades quickly. For example, increasing temperature from 0 to $1$ causes performance at high contamination levels ($1000$ replicas) to collapse by a factor of 40, down to the baseline performance of uncontaminated models. This suggests that contamination-driven performance is brittle; small adjustments in inference settings can eliminate the benefits of contamination almost entirely.
Figure 7: Solution Length and Temperature Together Govern How Quickly Memorization Decoheres. Top: Most models and contamination pairs exhibit scaling laws (Eqn. 3) as a function of the sequence length. For exceptions, see Appendix D. Bottom: The cumulative probability of generating a memorized solution can exist in one of three regimes: (1) Exponentially Fast Decoherence, (2) Brittle Memorization, and (3) Deterministic Lock-In. Sampling temperature can shift a model between regimes.
Finding #7: Longer Solution Length Mitigates Gains from Contamination
@@
To understand how solution length constrains performance, we binned problems into 10 log-spaced intervals ranging from the shortest (15 tokens) to the longest (1949 tokens). We found that Math Verify scores decrease significantly as solution length increases (Fig. 6). We identified a striking shift in the functional form of this decay based on contamination level: At higher contamination levels, the decay follows an approximate power law, but at lower levels of contamination, performance decays exponentially quickly to the noise floor. This suggests that maintaining a coherent memorized chain becomes increasingly difficult as the sequence grows, and also aligns with findings that longer sequences require more repetitions to be memorized (Jiang et al., 2025; Lu et al., 2024), though we demonstrate this here through controlled pretraining.
@@
Furthermore, solution length interacts with sampling temperature: For short solutions ($\leq 100$ tokens) at high contamination (316 replicas), raising the temperature from 0 to $1.0$ drops accuracy by $\sim 45\%$ on the largest model. However, for solutions of $400$ tokens, the same temperature increase causes accuracy to drop by nearly $100\%$.
Finding #8: Temperature and Solution Length Together Govern the Survival Process
Generative evaluations differ fundamentally from discriminative tasks because they require the model to maintain a coherent trajectory over a sequence of length $T$. For models that solve the task via memorizing solutions, we can characterize how the model must “survive” the risk of decoherence at every token step $t$ via studying the probabilities under teacher forcing.
To begin, motivated by prior work on in-context scaling laws (Anthropic, 2023; Team et al., 2024; Xiong et al., 2024; Anil et al., 2024), we find that for most model sizes and number of test set replicas, the negative log likelihood scales with the token index as a power law plus an irreducible error:
$\mathcal{L}_{t}(N,R)\approx E(N,R)+A(N,R)\cdot t^{-\alpha(N,R)}.$ (3)
However, we do find a small number of notable exceptions (Appendix. D). The probability of successfully generating a solution of length $T$ can then be approximated as:
$P(T)\approx\exp\left(-E(T-1)-\frac{A}{1-\alpha}(T^{1-\alpha}-1)\right).$ (4)
To explain intuitively, there are two opposing forces at the intersection of memorization and generation: The probability of generating a sequence decays geometrically (exponentially) with sequence length, making longer sequences harder to generate, but the the probability of knowing the correct next token grows geometrically with sequence length, making predicting the next token easier. These two rates reveal three distinct regimes of memorization (Fig. 8):
Regime I: Exponentially Fast Decoherence ($E>0$).
In the absence of meaningful contamination, the irreducible error term $E$ dominates. The survival probability decays exponentially ($P(T)\sim e^{-E\cdot T}$), rendering the generation of long solutions statistically improbable.
Regime II: Brittle Memorization ($E\approx 0,\alpha\leq 1$).
At intermediate contamination, the probability follows a Weibull-like stretched exponential ($P(T)\sim e^{-kT^{1-\alpha}}$). While the model can accurately generate sequences, its grasp on the solution is fragile, and the cumulative probability of success eventually vanishes.
Regime III: Deterministic Lock-In ($E\approx 0,\alpha>1$).
At high contamination, we observe a critical phase transition: $\alpha>1$ flips the sign of the exponent in Eq. 4, causing the penalty term $T^{1-\alpha}$ to decay to zero. Consequently, $P(T)$ converges to a non-zero constant even as $T\to\infty$. In this regime, the model’s certainty increases faster than the sequence accumulates risk, enabling the deterministic lock-in of long solutions.
The Effect of Sampling Temperature.
While the native scaling $\alpha$ is a property of the trained model, inference behavior is heavily modulated by the sampling temperature $\tau$. The dependence arises from the mechanics of the softmax distribution: in the regime of memorization, the model’s unnormalized confidence (the logit gap) grows logarithmically with context length ($\Delta z_{t}\propto\alpha\ln t$). Since temperature acts as a scalar divisor on logits ($z\to z/\tau$), it modifies the rate at which probability mass concentrates on the memorized path. This yields an effective scaling exponent:
$\alpha_{\mathrm{eff}}(\tau)\approx\alpha/\tau.$ (5)
This relationship reveals that temperature can artificially shift a model between regimes. Low-temperature sampling ($\tau<1$) inflates the exponent ($\alpha_{\mathrm{eff}}>\alpha$), potentially pushing a model from Brittle Memorization into pseudo-stable Deterministic Lock-In. This exposes the fragility of memorization: a model operating in Deterministic Lock-In ($\alpha_{\mathrm{eff}}>1$) at greedy settings ($\tau\to 0$) can be instantaneously reverted to Brittle Memorization ($\alpha_{\mathrm{eff}}<1$) simply by increasing $\tau$. For example, if a model has a natural scaling $\alpha=0.8$, greedy decoding effectively amplifies this to $\alpha_{\mathrm{eff}}\gg 1$, feigning robust knowledge. However, sampling at $\tau=1.0$ exposes the true exponent $\alpha<1$, causing failure for long sequences. This mechanism explains why high-temperature sampling acts as a “truth serum,” decoupling the gains of contamination from robust generalization.
Figure 8: Regimes of Memorization: Brittle vs. Lock-In. Phase diagrams illustrating the maximum sustainable solution length $T$ (hue) for a fixed survival probability $P(T)=0.01$, plotted against the irreducible error $E$ and scaling coefficient AA.
@@ 6 Discussion
Our application of neural scaling laws reveals the economic magnitude of this threat. Under standard scaling law assumptions, we demonstrate that a single replica of the test set allows a model to bypass the estimated “irreducible error” barrier, effectively simulating performance that would otherwise require an infinite amount of compute on uncontaminated data. However, this performance is brittle. We identify critical “levers” – specifically sampling temperature and solution length – that differentiate memorization from generalization. Unlike robust reasoning, which survives stochastic sampling, contamination-driven performance collapses under high temperatures. Similarly, the shift from power law to exponential decay in performance as solution length increases signals the limits of the model’s ability to hold a memorized chain without decoherence.
Perhaps most critically for practitioners, standard training pipelines can mask or expose this issue in counter-intuitive ways. The interaction between contamination and further training is governed by a “dose-response” effect (Schaeffer et al., 2025c): overtraining on fresh data dilutes the contamination, and Supervised Fine-Tuning (SFT) on valid training data can actually degrade test performance by overwriting memorized test samples. This implies that a drop in test accuracy after SFT, usually a sign of alignment tax or forgetting, may actually be a positive signal of decontamination.
@@
We focused on a single generative benchmark, MATH, to enable automatic verification and controlled contamination. Consequently, our findings may not fully capture how contamination behaves for other tasks such as coding or creative writing. Additionally, our experiments utilized decoder-only dense transformer models (Qwen 3) up to 344M parameters; results at this scale may not extrapolate to larger models or other architectures. Finally, our pretraining corpus represents a specific mixture of web-crawl data; specialized or heavily filtered corpora could alter the base difficulty of memorization and the specific thresholds at which contamination effects become visible.
@@
Several avenues for future research:
  • Reconciling Memorization Thresholds: Our discovery that a single test set replica drives loss below the irreducible error contrasts with recent assertions that single-shot verbatim memorization is an “illusion” (Huang et al., 2024) or that membership inference is fundamentally capped (Hayes et al., 2025). Future work should investigate what the explanatory differences are; one possibility is that memorization depends on the amount relative to corpus size (Schaeffer et al., 2025c; Wei et al., 2025).
  • Architectural Susceptibility: We studied dense models, but it remains an open question whether Mixture-of-Experts (MoEs) or State Space Models (SSMs) exhibit different memorization patterns.
  • Harder Benchmarks: We hypothesize that contamination efficacy is inversely correlated with problem compressibility. Investigating whether harder benchmarks (e.g., AIME) require higher “doses” of leakage to achieve the same performance boost would be a valuable contribution to benchmark design.
@@
An earlier version of this manuscript was peer reviewed and presented at the NeurIPS 2025 Workshop on Evaluating the Evolving LLM Lifecycle: Benchmarks, Emergent Abilities, and Scaling (Schaeffer et al., 2025d).
@@ 7 Acknowledgments
RS acknowledges support from Stanford Data Science and from the OpenAI Superalignment Fast Grant. JK acknowledges support from NSF grant number DGE1656518. SK acknowledges support by NSF 2046795 and 2205329, the MacArthur Foundation, Stanford HAI, OpenAI and Google Inc. KZL is generously supported by the Amazon AI PhD Fellowship. AA is generously supported by a Knight-Hennessy Fellowship, an NSF Graduate Research Fellowship, and a Georgetown Foundation Research Grant. Table 1 experiments were run using a computing grant provided by TensorPool.
@@ Appendix B Pretraining Implementation Details
Model Architecture
We pretrained Qwen 3 (Yang et al., 2025a) architecture causal language models from random initialization. Table 2 shows the depth (number of layers) and width (hidden size) configurations for each model size. The intermediate size for the feed-forward layers follows Qwen 3’s formula: $256\cdot\lfloor(255+\lfloor 8\cdot\text{hidden\_size}/3\rfloor)/256\rfloor$. All models used Flash Attention 2 (Dao, 2023) and bfloat16 precision.
Table 2: Model architecture configurations following Qwen 3 scaling patterns.
Parameters Num. Layers Hidden Size
34M 3 96
62M 5 160
93M 6 224
153M 9 320
344M 14 576
Optimizer and Learning Rate
We used the AdamW optimizer (Loshchilov & Hutter, 2019) with HuggingFace defaults (i.e., $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$, weight decay =0=0). We used linear warmup for 250 steps followed by cosine annealing to zero. Following Shuai et al. (2024), we scaled the batch size with the total number of training tokens $D$ as:
$\text{tokens per optimizer step}=3.24\times 10^{3}\times D^{0.264}.$ (6)
The learning rate was scaled with the square root of the batch size: $\eta=10^{-6}\times\sqrt{\text{tokens per optimizer step}}$. Gradients were clipped to a maximum norm of 1.0.
Data Mixing and Contamination
[…]
Distributed Training
Training used PyTorch’s DistributedDataParallel (DDP) with the NCCL backend. All experiments were logged to Weights & Biases, and trained models were uploaded to HuggingFace Hub.
For more information, please see our public GitHub repository.
Appendix C Math Verify Scoring Bug
During our experiments, we discovered that EleutherAI’s Language Model Evaluation Harness (Gao et al., 2024) contained a bug in its Math Verify scoring implementation that caused systematically incorrect results. This bug affected the minerva_math task and related mathematics evaluation tasks.
The Problem
The minerva_math task’s utility code called remove_boxed() on extracted answers before passing them to math_verify.parse(). The original implementation was:
res = verify(parse(doc["answer"]), parse(candidates))
[…]
This caused Math Verify scores to return 0 even when the model’s answer was mathematically correct. As a diagnostic, we evaluated the benchmark’s own gold reference solutions and observed Math Verify scores of approximately 70%—a clear indication that the scoring mechanism itself was flawed rather than the solutions.
The Fix
We reported this issue and a fix was merged that passes the full solution text to math_verify.parse(), allowing it to use its built-in extraction heuristics:
res = verify(gold=parse(doc["solution"]), target=parse(candidates))
[…]
Implications
This bug was present in versions of the evaluation harness prior to v0.4.8 (August 2025). Any research reporting MATH benchmark scores using Math Verify from affected versions may have systematically underestimated model performance. We recommend that researchers using these benchmarks verify they are using a corrected version of the evaluation code.
Appendix D Exceptions to In-Context / Sequence Scaling Laws
In Sec. 5, we reported that most model sizes and number of test set replicas exhibit in-context scaling laws (Eq. 3). While the power law plus irreducible error model provides an excellent fit for the majority of experimental conditions, we observe a systematic deviation in specific edge cases. In these instances, the negative log-likelihood (NLL) increases with token index $t$ rather than decaying. We identified significant positive slopes in the NLL-vs-token trajectory for two distinct groups of models.
Group I: Capacity Limitations (Uncontaminated Regime)
For the smallest uncontaminated model (34M, $R=0$), the NLL increases monotonically from token 1 to token 645. This behavior is consistent with limited effective context windows in small architectures. Without the aid of memorization ($R=0$), the 34M parameter model struggles to model long-range dependencies, causing prediction quality to degrade as the necessary context grows beyond its effective capacity. Larger uncontaminated models (62M+) do not exhibit this behavior, showing standard NLL decay.
Group II: Selection Bias (High Contamination Regime)
The most pronounced uptick occurs in the largest, most contaminated models (e.g., 344M, $R=3162$), where NLL is extremely low ($\sim 10^{-3}$) but rises sharply at late token indices. Our analysis suggests this is driven by selection bias inherent to the dataset structure.
Because the MATH dataset contains solutions of varying lengths, the set of problems contributing to the NLL at $t=800$ is a small, non-random subset of the problems at $t=100$.
  • Survivor Bias: Only the 100 longest solutions in the test set reach token 800.
  • Differential Memorizability: We observe a strong correlation between solution length and residual NLL in the high-contamination regime. The uptick indicates that these long surviving sequences are systematically “harder” to memorize than the shorter sequences that drop out earlier.
We verified that these deviations are not artifacts of data corruption. The token-level sample counts are consistent across all runs, with $100\%$ of sequences (N=5,000) present at token 0, dropping to $2\%$ (N=100) at token 800. This attrition is strictly deterministic, governed by the length distribution of the MATH dataset.
Figure 9: Deviations from In-Context / Sequence Scaling Laws. Two groups of models exhibit increasing negative log likelihoods with increasing token index: smaller uncontaminated models and larger massively contaminated models.
Figure 10: Fit Parameters for In-Context / Sequential Scaling Laws by Model Size and Number of Test Set Replicas.
Figure 11: Math Verify Score Is Correlated with Pretraining Loss. Math Verify scores correlate strongly with the cross entropy loss achieved on the MATH Test Set during training, where differences in these graphs are attributable to increased repetition on the benchmark test set. The correlation is significantly weaker for high temperatures and falls to nearly 0 for temperatures above $1.0$.